# Critical behavior and multistability in quasi-two-dimensional turbulence

### v1

@@ Abstract
Two-dimensional (2D) turbulence, despite being an idealization of real flows, is of fundamental interest as a model of the spontaneous emergence of order from chaotic flows. The emergence of order often displays critical behavior, whose study is hindered by the long spatial and temporal scales involved. Here, we experimentally study turbulence in periodically driven nanofluidic channels with a high aspect ratio using superfluid helium. We find a multistable transition behavior resulting from cascading bifurcations of large-scale vorticity and critical behavior at the transition to quasi-2D turbulence consistent with phase transitions in periodically driven many-body systems. We demonstrate that quasi-2D turbulent systems can undergo an abrupt change in response to a small change in a control parameter, consistent with predictions for large-scale atmospheric or oceanic flows.
Chaotic, turbulent flows are ubiquitous and occur in vastly diverse systems ranging from a flow around a swimming penguin Masud et al. (2022) to climate systems on Earth Boffetta and Ecke (2012). In homogeneous isotropic turbulence in three dimensions, kinetic energy is transported to progressively smaller scales until dissipated by viscosity, thus destroying any large-scale features Davidson (2015). However, confined to two dimensions, turbulence transports energy to large scales in an inverse cascade, which can result in large-scale self-organization Boffetta and Ecke (2012).
The self-organized state is not necessarily unique, and sudden changes between distinct metastable states are possible. Bifurcations in quasi-2D flows are of particular relevance to the dynamics of geophysical flows. For example, large-scale circulating flow in the Atlantic Ocean is likely to transition into a new regime with a significantly altered global climate Ditlevsen and Ditlevsen (2023); van Westen et al. (2024). More detailed simulations suggest a rich landscape of multiple stable states of large geophysical flows Lohmann et al. (2024). Numerical models of idealized 2D turbulence similarly show multiple stable states, such as large vortices and zonal jets between which the flow transitions stochastically Bouchet and Simonnet (2009), although these transitions are rare Bouchet et al. (2019). There is a growing body of evidence, however, that turbulent flows display analogous behavior to phase transitions in condensed matter systems Cortet et al. (2010) or various nonequilibrium phase transitions (e.g., directed percolation) Sano and Tamai (2016); Hof (2023).
Critical behavior in quasi-2D flows is unexpectedly rich and remains poorly understood Alexakis (2023). Therefore, in this work, we study driven oscillatory turbulent flow confined to channels of aspect ratio $2000D\times 2000D\times D$, $D\approx 500$ nm. We experimentally show critical behavior consistent with random organization Corté et al. (2008) and that the turbulence develops through a cascade of saddle-node bifurcations when the large-scale friction is sufficiently low, and the forcing length scale is sufficiently small.
@@
Flows of superfluid 4He are a superposition of viscous normal flow and inviscid superflow Tilley and Tilley (1990). Microscopically, the superfluid flow is irrotational Tilley and Tilley (1990). Nonzero circulation can exist around topological defects in the superfluid, quantized vortices, which carry a single quantum of circulation $\kappa\approx 9.97\times 10^{-8}$ m2s-1. Quantized vortices facilitate friction-like coupling (mutual friction) of superfluid and normal fluid flows Barenghi et al. (1983). On scales larger than the intervortex spacing, the flow of superfluid helium can be described using coupled equations for superflow and normal fluid flow Donnelly (1991), which in our case reduce to $\mathbf{v}_{n}\approx 0$ (due to viscous clamping Rojas and Davis (2015); Souris et al. (2017)) for normal velocity and for superfluid velocity $\mathbf{v}_{s}$ Donnelly (1991)
@@
We study two-dimensional flows using nanofluidic Helmholtz resonators, shown in Fig. 1(a), which consist of a circular volume (“basin”) connected to the surrounding helium bath at saturated vapor pressure by two channels. Similar devices have previously been used for the study of 3He superfluidity Shook et al. (2020, 2024), finite-size effects in 4He Varga et al. (2022), quasi-two-dimensional turbulence Varga et al. (2020) and superfluid optomechanics Spence et al. (2023). The superfluid helium in the channels oscillates under the restoring force provided by the compressibility of the helium and the deformation of the quartz substrate Rojas and Davis (2015). The top and bottom walls of the nanofluidic volume are coated with a 50 nm aluminum layer used for capacitive actuation and sensing (see appendix Supplementary Information and Rojas and Davis (2015); Souris et al. (2017) for details). We studied three flow geometries with different energy injection length scales, whose velocity profile of the Helmholtz mode is shown in Fig. 1(a) — a flow around a sharp corner (device C); a constriction producing an oscillating jet (device J) and a flow through a regular grid (device G; see appendix Supplementary Information for exact dimensions and data from additional device similar to C with rounded corners). Height $D\approx 500$ nm was used for all geometries.
@@
The high aspect ratio of the flow indicates that the turbulence will be two-dimensional Benavides and Alexakis (2017); Ecke (2017). However, in superfluid helium, the core of the quantized vortex sets the smallest length scale, approximately 0.15 nm Tilley and Tilley (1990). Numerical simulations based on the vortex filament model (see appendix Supplementary Information Sec. 6.2) show that the lifetime of intrinsically three-dimensional vortex configurations is significantly shorter than the flow oscillation period. They are thus transient and cannot participate in long-lived turbulent structures.
Figure 1: Nanofluidic devices and their response. (a) Typical dimensions of the nanofluidic Helmholtz resonators. Dark-shaded areas show aluminum electrodes deposited on the top and bottom walls. The 0.5 mm quartz substrates are not shown. The entire device is submerged in superfluid 4He. A finite element simulation shows the velocity profile of the used acoustic mode. (b) Typical resonant response of the device normalized by the drive. (c) Peak velocity as a function of the driving amplitude of device C at 1.45 K, 1.65 K, and 1.85 K (top to bottom); shaded areas indicated the excess nonlinear damping.
To observe the development of turbulence, the Helmholtz resonance was excited on its resonance frequency with slowly swept drive amplitude in increasing and decreasing directions. The typical response showing the excess dissipation past a critical velocity is shown in Fig. 1(c). Since the flow can transition stochastically between several states, the drive sweep is repeated in several cycles to fully sample the possible states. The vortex density calculated from the excess damping (appendix Supplementary Information (S17)) is shown in Fig. 2(a–i) as two-dimensional histograms for three geometries for a high, intermediate, and low temperature (1.85 K, 1.65 K and 1.45 K, respectively), which determines the large-scale friction in (1). Data from all devices and temperatures are shown in appendix Supplementary Information, Figs. S11–S14.
@@
At high temperatures above a critical velocity, the observed vortex density gradually and reversibly increases above zero for all geometries. For lower temperatures, the behavior strongly depends on geometry. For device C (panels (a–c)), weak hysteresis is observed after the first critical velocity. For device J (panels (d–f)), the flow transitions to a regime with increased $L$ through a discontinuity with stochastic switching between two discrete states. For device G (panels (g–i)), we observe several bifurcations below approximately 1.65 K, resulting in multiple metastable states and hysteresis in the transition region. Additionally, at least two intermittently switching transitions are observable in panel (g). In Fig. 2(j), we show the bifurcation diagram of the branching of the vortex density as a function of temperature. Each branch separated by discontinuous transitions from the others is characterized by a dimensionless quantity $\langle L\kappa^{2}/v^{2}\rangle$. For device J, two branches are observed from 1.85 K to 1.4 K, where the upper branch divides into two. For device G, the branches proliferate in a fractal fashion below 1.45 K. However, the growth in complexity results in a less reliable classification of the data into separate branches.
@@
Figure 2: Vortex density bifurcation diagrams. (a–i) Two-dimensional histograms of vortex density and velocity (darker color means higher probability; logarithmic color scale) for device type C (panels a–c), J (d–f) and G (g–i) for three temperatures, 1.45 K (a, d, g), 1.65 K (b, e, h), and 1.85 K (c, f, i). (j) The temperature-density bifurcation diagram shows the average value of $L\kappa^{2}/v^{2}$ calculated for each individual branch of $L(v)$. Red, device J, black, device G. For $T<1.45$ K, the data for device G continue to branch in a fractal manner (see appendix Supplementary Information, Fig. S13).
To describe the appearance of the discontinuous transitions in devices J and G, we seek a low-dimensional model of the vortex density $L$ (similarly to ref. Varga et al. (2020)). The behavior can be qualitatively captured with a dynamical system (for dimensionless quantities)
@@
where $L_{j}$ is the vortex density in mode $j$ (i.e., a wave vector for a spatially infinite or periodic system), $j=1...N$, where $N$ is the number of participating modes; $d_{l}$ represents the strength of the large-scale damping due to mutual friction (first term) relative to the pair annihilation of colliding vortices (second term); and $g_{j}$ represents the forcing, which, in general, will be a nonlinear function of the flow velocity and geometry. For the total vortex density, we take $L^{2}=\sum_{j}L^{2}_{j}$. The essential feature of the model is that, for large populations $L>1$, the annihilation of vortices is suppressed. A typical curve $\mathrm{d}L/\mathrm{d}t$ for $N=1$ is shown in Fig. 3(a), and a bifurcation diagram with drive $g$ as the control parameter is shown in Fig. 3(b). For low damping $d_{l}$, the time derivative exhibits a local minimum, resulting in three fixed points, two stable and one unstable, for a certain range of $g$, which qualitatively describes the bifurcation seen in device J. With increasing temperature, the mutual friction increases, increasing large-scale friction on the vortex system. Indeed, the bistability is lost when the parameter $d_{l}$ is increased, similarly to the experiment (cf. Fig. 3(b) and Fig. 2 middle row).
@@
The energy injection scale $l_{f}$ for the device J is close to the width of the constriction, about half of the maximum size of the system $l_{0}$. Therefore, only a small number of wave vectors can exist in the inverse cascade between $l_{f}$ and $l_{0}$. In device G, we increase the ratio $l_{0}/l_{f}$, thus extending the number of modes participating in the inverse cascade. Therefore, we set $N>1$ and assume that the forcing is localized at $j=1$, $g_{1}=g$ and that the higher modes are forced in a cascading manner $g_{j}=cL_{j-1}$, $j>1$, where $c$ characterizes the mode coupling ($d_{l}$ and $c$ are kept constant for simplicity). For $d_{l}=0.05$, $c=0.04$, the total density $L$ in Fig. 3(c) shows multistability due to overlapping pairs of saddle-node bifurcations resembling the experimental data of device G (full lines show stable stationary solutions; dashed lines the unstable fixed points). Although the coupling between vorticity modes is simplified Kraichnan and Montgomery (1980), multistable behavior can occur for more complicated multiplicative coupling Huang et al. (2016). Extending the interpretation of the single-mode dynamics of device J, the resulting physical picture is that, as a particular mode is sufficiently excited, the vortices cluster in a certain pattern to reduce pair annihilation. The energy in the mode thus increases, which allows nonlinear interactions to force progressively larger scales until the size of the system is reached.
The suppression of the annihilation is responsible for the appearance of a pair of bifurcations and the bistable regime. The exact form of the annihilation term does not affect the qualitative behavior of the solutions as long as a local minimum appears. The rational annihilation term in (2) was chosen for simplicity and to maintain the expected $L^{2}$ behavior for low $L$ corresponding to the annihilation of randomly distributed vortices Varga and Skrbek (2018). Since the vortices are expected to be approximately point-like (see appendix Supplementary Information Fig. S7), suppression of annihilation at high densities can occur due to spatial clustering of like-signed vortices, suggesting the emergence of order.
@@
Note that the bistability observed here is different than observed previously in 1 µm confinement Varga et al. (2020), where the bistability was long-lived on the scale of the entire duration of the experiment (i.e., minutes), whereas here we observe intermittent switching or pronounced multistability. However, present results continue the trend observed in Varga et al. (2020) of weakened bistability for C-type device as $D$ decreases. Both types of multi-stability should be considered as special cases of more general dynamical behavior, a complete understanding of which is not yet available.

### v2

@@ Abstract
Two-dimensional (2D) turbulence, despite being an idealization of real flows, is of fundamental interest as a model of the spontaneous emergence of order from chaotic flows. The emergence of order often displays critical behavior with the involvement of a broad range of spatial and temporal scales. Here, we experimentally study turbulence in periodically driven nanofluidic channels with a high aspect ratio using superfluid helium. We find a multistable transition behavior resulting from cascading bifurcations of large-scale vorticity and critical behavior at the transition to quasi-2D turbulence consistent with phase transitions in periodically driven many-body systems. We demonstrate that quasi-2D turbulent systems can undergo an abrupt change in response to a small change in a control parameter, consistent with predictions for large-scale atmospheric or oceanic flows.
I Introduction
Chaotic, turbulent flows are ubiquitous and occur in vastly diverse systems ranging from a flow around a swimming penguin [1] to climate systems [2, 3]. In homogeneous isotropic turbulence in three dimensions, energy is transported to small scales until dissipated by viscosity, thus destroying any large-scale features [4]. However, in two dimensions, turbulence transports energy to large scales in an inverse cascade, which can result in the emergence of large-scale coherent structures called condensates in analogy with Bose-Einstein condensation [2]. The condensate state is, generally, not unique, and stochastic changes between distinct metastable states, such as zonal jets or large-scale vortices, are possible [5]. These transitions are rare [6]; however, recent extensive simulations [7] show a transition between preferred states as Reynolds number and flow aspect ratio is varied with fluctuating transitions in the intermediate region. Allowing for finite thickness of the flow, the possible phase transition-like phenomena in turbulent flows become very rich [8]. However, even the existence of a critical point for the lifetime or the formation time of the vorticity condensate is uncertain, where scenarios with zero, one or two critical points are possible [9].
Experimentally, phase transitions were observed in 3D turbulent flows [10], and multistability of turbulent flows was observed as spontaneous reversals in magnetically forced turbulence [11, 12] in thick layers, and for flows in the stratosphere [13] and their laboratory models [14]. In thin layer flows, experiments focused mostly on the build-up and long-time behavior of vorticity condensates and their influence on velocity statistics in the inverse cascade [15, 16]. Long-lived bistability was observed in oscillatory flows [17], although no stochastic transitions were present.
The experimental study of intermittent transitions is often limited by the aspect ratios required for two-dimensionality and the rarity of the transitions. In this work, we aim to shed light on the critical behavior of quasi-two-dimensional turbulence, which is thus far accessible only in numerical simulations. Our experimental method allows long observation times compared to typical timescales of the turbulent flow. We observe the stochastic transitions between bistable states and strong multistability involving at least 8 turbulent states in driven oscillatory turbulent flow of superfluid helium confined to channels of aspect ratio $2000D\times 2000D\times D$, $D\approx 500$ nm. The lifetimes of the turbulent states scale with the flow velocity in the bistable regime approximately as a power-law with critical exponent consistent with random organization [18] or 2D directed percolation [19], which supports the scenarios with at least one critical point for the formation of the condensate suggested in [9].
The paper is organized as follows: In section II, the Helmholtz resonators and the measurement of vortex line density are briefly reviewed; in section III.1, we discuss the first critical velocity for the onset of nonlinear damping, which is likely connected with depinning of quantized vortices from the surface; in section III.2 we show the growth of vortex line density with superflow velocity in the transitional region and the appearance of multistability depending on the geometry and temperature and develop a dynamical model that qualitatively captures the observed behavior; in section III.3 the stochastic transitions between two different flow states are shown whose lifetimes approximately follow diverging power laws after which the discussion and conclusions follow.
II Experimental Setup
@@
We use superfluid 4He (He-II) as the working medium, whose lack of viscosity allows for extremely narrow flow channels. Flows of He-II are a superposition of viscous normal flow and inviscid superflow [20]. Microscopically, the superfluid flow is irrotational. However, circulation can exist around topological defects in the superfluid, quantized vortices, which carry a single quantum of circulation $\kappa\approx 9.97\times 10^{-8}$ m2s-1 [20]. Quantized vortices facilitate friction-like coupling (mutual friction) of superfluid and normal fluid flows [21]. On scales larger than the intervortex spacing, the flow of superfluid helium can be described using coupled equations for superflow and normal fluid flow [22], which in our case reduce to $\mathbf{v}_{n}\approx 0$ due to viscous clamping [23, 24] for normal velocity and for superfluid velocity $\mathbf{v}_{s}$ [22]
@@
We study two-dimensional flows using nanofluidic Helmholtz resonators, shown in Fig. 1(a), which consist of a circular volume (“basin”) connected to the surrounding helium bath at saturated vapor pressure by two channels. Similar devices have previously been used for the study of 3He superfluidity [28, 29], finite-size effects in 4He [30], quasi-two-dimensional turbulence [17] and superfluid optomechanics [31, 32]. The superfluid helium in the channels oscillates under the restoring force provided by the compressibility of the helium and the deformation of the quartz substrate, resulting in a Helmholtz resonance [23]. The top and bottom walls of the nanofluidic volume are coated with a 50 nm aluminum layer used for capacitive actuation and sensing (see Appendix A for details on the detection circuit). A semi-open nanofluidic channel is etched onto a fused silica substrate, two of which are then bonded together to form a fully enclosed nanofluidic cavity approximately 500 nm in height (see [23, 24] for further fabrication details). We studied four flow geometries with different energy injection length scales, whose velocity profile of the Helmholtz mode is shown in Fig. 1(b,c) — a flow around a sharp corner (device C); a similar device with rounded-off corners (device R); a constriction producing an oscillating jet (device J) and a flow through a regular grid (device G). In device R, the behavior was similar to device C; however, reliable velocity calibration was not possible. The data obtained from this device is available only in the attached dataset. Height $D\approx 500$ nm was used for all geometries. The characteristic length scale at which a viscous fluid follows the motion of an oscillating wall is the viscous penetration depth $\delta=\sqrt{2\nu/\omega}$, where $\nu$ is the kinematic viscosity and $\omega$ the flow oscillation frequency. For the typical resonance frequency 2 kHz and temperature 1.65 K (see Fig. 8 for the temperature dependence of the resonance frequency and linewidth), and using the normal fluid kinematic viscosity from [25], $\delta\approx$2.7 $\mu$m, which is 5-times greater than the spacing between the upper and lower walls of the cavity. We can, therefore, consider the viscous normal fluid component of superfluid helium to be at rest.
The typical flow velocity amplitude as a function of the drive is shown in Fig. 2(b,c). For low velocities, the scaling with the applied force remains linear, determined primarily by thermoviscous dissipation (see Appendix C and [24]). Past a critical velocity, nonlinear damping sets in. The high aspect ratio of the flow indicates that the turbulence will be two-dimensional [33, 34]. However, the quantized vortex sets the smallest length scale, approximately 0.15 nm [20]. Thus, in principle, vortex loops or half-loops attached to the walls created, e.g., during reconnections, might exist in the flow. However, numerical simulations based on the vortex filament model (see Appendix D) show that the lifetime of three-dimensional vortex configurations is significantly shorter than the flow oscillation period (several tens of microseconds, depending on the specific geometry of the loop and its orientation with respect to the flow). They are thus transient and cannot participate in long-lived turbulent structures.
Figure 1: Nanofluidic devices. (a) Typical dimensions of the nanofluidic Helmholtz resonators. Dark-shaded areas show aluminum electrodes deposited on the top and bottom walls. The 0.5 mm quartz substrates are not shown. The entire device is submerged in superfluid 4He. (b) Details of the geometry of the four resonator channels used labeled C, R, J, and G (left to right) in the text. The height of all four cavities is 500 nm. The aluminum electrodes are not shown. (c) A finite element simulation shows the velocity profile of the used acoustic mode.
Figure 2: Nonlinear response of the Helmholtz resonators. (a) Mutual friction coefficients entering (1) from ref. [25]. (b) Typical resonant response of the device normalized by the drive. Suppression of the peak indicates nonlinear dissipation. (c) Peak velocity as a function of the driving amplitude of device C at 1.45 K, 1.65 K, and 1.85 K (top to bottom); shaded areas indicated the excess nonlinear damping.
II.1 Measurement of vortex line density
Similarly to the second sound [35], the fourth sound used in this work is damped by the presence of quantized vortices. The excess nonlinear damping (shaded areas in Fig. 2(c)) can be used to calculate the average number of vortices per unit area (vortex density $L$). For completeness, we briefly reproduce the derivation of the equation of motion of the Helmholtz mode and the damping due to quantized vortices based on refs. [17, 36].
$\delta\rho=\delta\left(\dfrac{M_{B}}{V_{B}}\right)=\dfrac{\delta M_{B}}{V_{B}}-\dfrac{M_{B}}{V_{B}^{2}}\delta V_{B}=\dfrac{1}{V_{B}}(-2a\rho_{s}y+2\rho Ax),$ (2)
$\delta P=\dfrac{1}{\rho\chi V_{B}}(2\rho Ax-2a\rho_{s}y).$ (3)
$F=k_{p}x+A\delta P,$ (4)
where $k_{p}$ is the spring constant, and we neglect the inertia of the basin walls. Combining (2), (3), and (4), we can write the equation of motion for the fluid displacement $y$ in the channel with length $l$
$\rho_{s}al\ddot{y}$ $=\dfrac{\rho_{s}}{\rho}a\delta P-F_{f},$ (5)
$\rho_{s}al\ddot{y}$ $=\dfrac{\rho_{s}ak_{p}}{\rho^{2}(\chi V_{B}k_{p}+2A^{2})}\left[\dfrac{2A\rho}{k_{p}}F-2a\rho_{s}y\right]-F_{f},$ (6)
where we added an additional friction force $F_{f}$. Here the driving force $F=C_{0}U_{0}(t)U_{B}/D$, since the total voltage is a sum of the bias voltage $U_{B}$ and the driving alternating voltage $U_{0}(t)$ and we only take the component that is resonant with the Helmholtz mode. The friction force $F_{f}$ can be divided into two contributions $F_{f}=F_{\mathrm{in}}+F_{\mathrm{ns}}$. The first part is the intrinsic damping of the superfluid motion [24], which is discussed in Appendix C, and can be written in the form
$F_{\mathrm{in}}=al\rho_{s}\gamma_{0}\dot{y},$ (7)
where $\gamma_{0}$ is a dimensionless intrinsic damping factor. The second contribution is the mutual friction force between the superfluid and quantized vortices, which can be expressed [22] as
$\mathbf{F}_{\mathrm{ns}}=al\alpha\rho_{s}\hat{\bm{\omega}}_{s}\times[\bm{\omega}_{s}\times(\mathbf{v_{\mathrm{n}}}-\mathbf{v_{\mathrm{s}}})]+al\alpha^{\prime}\rho_{s}\bm{\omega}_{s}\times(\mathbf{v_{\mathrm{n}}}-\mathbf{v_{\mathrm{s}}}),$ (8)
where $\bm{\omega}_{s}$ is the superfluid vorticity and $\hat{\bm{\omega}}_{s}$ is the unit vector with the same direction. Due to geometrical constraints we have for the vortex density $L\approx|\bm{\omega}_{s}|/\kappa$, where $\kappa\approx 9.97\times$10-8 m2s-1 is the quantum of circulation. Assuming that the normal fluid component is viscously clamped to the walls, i.e. $\mathbf{v_{\mathrm{n}}}=0$, and neglecting the second non-dissipative mutual friction term, (8) simplifies to
$F_{ns}=al\alpha\kappa\rho_{s}Lv_{\mathrm{s}}\sin^{2}{\theta},$ (9)
where the angle $\theta$ is between $\mathbf{v_{\mathrm{s}}}$ and $\hat{\bm{\omega}}_{s}$. Assuming that the vortices are preferentially oriented perpendicularly to the superfluid flow (a consequence of 2-dimensionality), we have $\sin^{2}{\theta}=1$.
The dynamics of $L(t)$ likely occur on a significantly longer time scale than the flow oscillation period, as evidenced by the relatively long-lived states in both the multi-stable and intermittently switching regimes (from roughly the inverse linewidth up to several seconds; Sec. III.2). In the steady state, we therefore consider $L$ to be a constant. This constant will, however, depend on the RMS flow velocity (and thus drive amplitude and frequency). Therefore, a Lorentzian peak shape of the resonant response should not be expected. A similar constancy of $L$ generated by the second sound in a 3D geometry was reported in turbulence generated by the second sound [37].
Identifying $\dot{y}$ with $v_{\mathrm{s}}$ and substituting friction (7),(9) into the equation of motion (6) results in
$\ddot{y}+(\alpha\kappa L+\gamma_{0})\dot{y}+\dfrac{2\rho_{s}ak_{p}}{l\rho^{2}(\chi V_{B}k_{p}+2A^{2})}y=\dfrac{2AC_{0}U_{B}}{\rho l(\chi V_{B}k_{p}+2A^{2})h}U_{0}(t),$ (10)
$\ddot{y}+\tilde{\gamma}\dot{y}+\omega^{2}_{0}y=f_{0}e^{i\omega t},$ (11)
$\tilde{\gamma}$ $=\alpha\kappa L+\gamma_{0}=\gamma(L)+\gamma_{0},$ (12)
$\omega_{0}^{2}$ $=\dfrac{2\rho_{s}ak_{p}}{l\rho^{2}(\chi V_{B}k_{p}+2A^{2})},$ (13)
$f_{0}(t)$ $=\dfrac{2AC_{0}U_{B}}{\rho l(\chi V_{B}k_{p}+2A^{2})h}U_{0}(t)=f_{0}e^{i\omega t}.$ (14)
$\dfrac{y_{0}}{y}-1=\dfrac{\gamma(L)}{\gamma_{0}}.$ (15)
Using Eq. 9 (with $\theta=\pi/2$) and the relation between the damping factor $\gamma_{0}$ and the width $\Delta_{0}$ of the un-damped resonance curve, $\gamma_{0}=2\pi\Delta_{0}$, results in
$L=\dfrac{2\pi\Delta_{0}}{\alpha\kappa}\left(\dfrac{y_{0}}{y}-1\right).$ (16)
In the present work, $y_{0}$ is extrapolated from the linear regime of the force-velocity curves shown in Fig. 2(c).
III Results
To observe the development of turbulence, the Helmholtz resonance was excited on its resonance frequency with slowly swept drive amplitude in increasing and decreasing directions. The typical response showing the excess dissipation past a critical velocity is shown in Fig. 2(c). Since the flow can transition stochastically between several states, the drive sweep is repeated in several cycles to sample the possible states fully.
III.1 First critical velocity
As the flow velocity increases, the nonlinear damping first appears at approximately 0.1 – 0.2 m/s (see appendix B for calibration of the measured electrical quantities to forces and velocities). For concreteness, we define the critical velocity as the velocity where the vortex line density calculated using (16) first exceeds a threshold of $L_{c}=5\times 10^{8}$ m-2 with results shown in Fig. 3(a). The velocities are only weakly dependent on device geometry and temperature, and the transition is generally reversible (with the exception of device G at the lowest attained temperatures, see Fig. 4).
A possible geometry and temperature-independent mechanism for the appearance of critical velocity is vortex pinning. Due to the small size of the vortex core, $a_{0}\approx 0.15$ nm, vortices will be pinned on surface roughness and will be unable to slide unless subject to superflow faster than the depinning velocity, which was estimated by Schwarz [38] for a semi-spherical protuberance of radius $b$ on a planar surface in a planar channel of height $D$ as
$v_{\mathrm{pin}}=\frac{\kappa}{2\pi D}\ln\left(\frac{b}{a_{0}}\right).$ (17)
Because of the stringent requirements of the surface cleanliness for the bonding of the resonators, any manipulation of the unbonded chips outside of the cleanroom environment is essentially destructive, therefore we were unable to measure the topography on the actual devices used in the study. However, a scan of the aluminum layer surface topography using atomic force microscopy is shown in Fig. 3(b) on devices created using a nearly identical fabrication process. The scan shows 10 µm square on the edge of the aluminum. The step height is 48.64 $\pm$ 0.04 nm, and the RMS surface roughness (that is, the root-mean-square distance to the mean plane) of the aluminum layer is approximately 4.1 nm.
Replacing the protuberance size $b$ in (17) with the measured surface roughness results for $D=500$ nm in $v_{\mathrm{pin}}\approx$ 11 cm/s, which is in good agreement with initial reversible transitions to nonlinear dissipation seen in Fig. 3(a). The origin of the slight temperature dependence of this critical velocity is unknown. However, (17) was obtained for steady flow for a single pinning site; in the present case, the temperature dependence is probably the result of the interaction of the temperature-dependent flow oscillation frequency and the vortex relaxation time given by the mutual friction [39].
@@ III.1 First critical velocity
Figure 3: First critical velocity. (a) The first critical velocity for the three device geometries. Within a factor of two, the velocities agree with the depinning velocity estimate (17). The critical velocity is determined as the velocity, where vortex line density increases above $5\times 10^{8}$ m-2 and the ends of the errorbars as velocities where $(5\pm 2.5)\times 10^{8}$ m-2 is exceeded. (b) Atomic force microscopy scan of the topography of the edge of the aluminum surface (higher, rougher area) in the basin area.
III.2 Hysteresis and multistability in turbulent transition
The vortex density calculated from the excess damping is shown in Fig. 4(a–i) as two-dimensional histograms for three geometries for a high, intermediate, and low temperature (1.85 K, 1.65 K and 1.45 K, respectively), which determines the large-scale friction in (1). Data from all devices and 14 temperatures in total are available in the attached dataset.
@@ III.2 Hysteresis and multistability in turbulent transition
At high temperatures above a critical velocity, the observed vortex density gradually and reversibly increases above zero for all geometries. For lower temperatures, the behavior strongly depends on geometry. For device C (panels (a–c)), weak hysteresis is observed after the first critical velocity. For device J (panels (d–f)), the flow transitions to a regime with increased $L$ through a discontinuity with stochastic transitions between two discrete states. For device G (panels (g–i)), we observe several bifurcations below approximately 1.65 K, resulting in multiple metastable states and hysteresis in the transition region. Additionally, at least two intermittent transitions are observable in panel (g). In Fig. 4(j), we show the bifurcation diagram of $L$ as a function of temperature. Each branch separated by discontinuous transitions from the others is characterized by a dimensionless quantity $\langle L\kappa^{2}/v^{2}\rangle$ (averaging over drive amplitudes within a given branch). For device J, two branches are observed from 1.85 K to 1.4 K, where the upper branch divides into two. For device G, the branches proliferate in a fractal fashion below 1.45 K. However, the growth in complexity results in a less reliable classification of the data.
@@
Figure 4: Vortex density bifurcation diagrams. (a–i) Two-dimensional histograms of vortex density and velocity (darker color means higher probability; logarithmic color scale) for device type C (panels a–c), J (d–f) and G (g–i) for three temperatures, 1.45 K (a, d, g), 1.65 K (b, e, h), and 1.85 K (c, f, i). (j) The temperature-density bifurcation diagram shows the average value of $L\kappa^{2}/v^{2}$ calculated for each individual branch of $L(v)$ (averaged over drives). Red, device J, black, device G. For $T<1.45$ K, the data for device G continue to branch in a fractal manner.
Two possibilities are currently available for the nature of discrete turbulent states: critical transition between forward and inverse cascades (e.g., [42]) or organization of the vorticity into condensates. Due to the existence of more than two discrete states, the latter option appears more likely. Thus, assuming that the vortex line density can cluster in several distinct distributions (“modes”, e.g., large-scale vortices or zonal jets [7]) we seek a low-dimensional model of the vortex density $L$ (similarly to ref. [17]) to describe the appearance of the discontinuous transitions in devices J and G. The behavior can be qualitatively captured with a dynamical system (for dimensionless quantities)
@@
where $L_{j}$ is the vortex density in mode $j$, $j=1...N$, $N$ is the number of participating modes, $d_{l}$ represents the strength of the large-scale damping due to mutual friction (first term) relative to the vortex pair annihilation (second term), and $g_{j}$ represents the forcing, which, in general, will be a nonlinear function of the flow velocity and geometry. For the total vortex density, we take $L^{2}=\sum_{j}L^{2}_{j}$. The essential feature of the model is that, for large populations $L>1$, the annihilation of vortices is suppressed. A typical curve $\mathrm{d}L/\mathrm{d}t$ for $N=1$ is shown in Fig. 6(a), and a bifurcation diagram with drive $g$ as the control parameter is shown in Fig. 6(b). For low damping $d_{l}$, the time derivative exhibits a local minimum, resulting in three fixed points, two stable and one unstable, for a certain range of $g$, which qualitatively describes the bifurcation seen in device J. With increasing temperature, the mutual friction increases, increasing large-scale friction on the vortex system. Indeed, the bistability is lost when the parameter $d_{l}$ is increased, similarly to the experiment (cf. Fig. 6(b) and Fig. 4 middle row).
@@
The energy injection scale $l_{f}$ for the device J is close to the width of the constriction, about half of the maximum size of the system $l_{0}$. Therefore, only a small number of wave vectors can exist in the inverse cascade between $l_{f}$ and $l_{0}$. In device G, we increase the ratio $l_{0}/l_{f}$, thus extending the number of modes participating in the inverse cascade. Therefore, we set $N>1$ and assume that the forcing is localized at $j=1$, $g_{1}=g$ and that the higher modes are forced in a cascading manner $g_{j}=cL_{j-1}$, $j>1$, where $c$ characterizes the mode coupling ($d_{l}$ and $c$ are kept constant for simplicity). For $d_{l}=0.05$, $c=0.04$, the total density $L$ in Fig. 6(c) shows multistability due to overlapping pairs of saddle-node bifurcations resembling the experimental data of device G (full lines show stable stationary solutions; dashed lines the unstable fixed points). Although the coupling between vorticity modes is simplified [43], multistable behavior can occur for more complicated multiplicative coupling [44]. Extending the interpretation of the single-mode dynamics of device J, the resulting physical picture is that, as a particular mode is sufficiently excited, the vortices cluster in a certain pattern to reduce pair annihilation. The energy in the mode thus increases, which allows nonlinear interactions to force progressively larger scales until the size of the system is reached. The suppression of the annihilation is responsible for the appearance of a pair of bifurcations and the bistable regime. The exact form of the annihilation term does not affect the qualitative behavior of the solutions as long as a local minimum appears. The rational annihilation term in (18) was chosen to maintain the expected $L^{2}$ behavior for low $L$ corresponding to the annihilation of randomly distributed vortices [45].
@@
Note that the bistability observed in device J is different than observed previously in 1 µm confinement [17], where it was long-lived on the time scale of the experiment (i.e., minutes). Here we observe intermittent switching or pronounced multistability. However, present results continue the trend observed in [17] of weakened bistability for C-type device as confinement $D$ decreases. Both types of multi-stability should be considered as special cases of more general dynamical behavior, a complete understanding of which is not yet available.